OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. c√c + √7x - 2c√c A.)11c B.)c√c C.)11c√c D.)11√c

OpenStudy (anonymous):

That should be c I think in place of x.. Right??

OpenStudy (anonymous):

Is this your question? \[c \sqrt{c} + \sqrt{7c} - 2 c \sqrt{c}\]

OpenStudy (anonymous):

wait no that's whats on my paper

OpenStudy (anonymous):

This should not be your question.. Check your question again and tell me what it is actually..

OpenStudy (anonymous):

my question is \[c \sqrt{c}+\sqrt{7x}-2c \sqrt{c}\] thats exactly what i have on the paper,

OpenStudy (anonymous):

But in no way, this will lead you to your answer as you are giving the answer choices..

OpenStudy (anonymous):

Check it once more, see you have x in your question but there is no x in the answer choices you are giving, something must be wrong..

OpenStudy (anonymous):

i agree with waterineyes

OpenStudy (anonymous):

Can you take screen shot or in anyway you can upload it here so that we can also see it..??

OpenStudy (anonymous):

thats EXACTLY what is says but they could have made a mistake, yah ill try give me a second

OpenStudy (anonymous):

Take your time, we are in no hurry..

OpenStudy (anonymous):

hey waterineyes are you familiar with tsa?

OpenStudy (anonymous):

tsa?? short form it is??

OpenStudy (anonymous):

thats what it says! lol they messed up lol i think they ment to put c

OpenStudy (anonymous):

the person who printed this must of been on coke :)

OpenStudy (anonymous):

where the heck did this nonsense come from??

OpenStudy (anonymous):

Wait.. What a question!! By putting c, you will not get your answer... The question is totally and completely wrong..

OpenStudy (anonymous):

even if the \(x\) was a \(c\) you till have no correct answers given

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

I think : \(\sqrt{7x}\) will be : \(12c \sqrt{c}\).. Ha ha ha ha..

OpenStudy (anonymous):

they did make a mistake lol that x should have been a "c" lol my answer i got was \[11c \sqrt{c}\] (:

OpenStudy (anonymous):

and i was right (:

OpenStudy (anonymous):

How you got 11c root(c) ??

OpenStudy (anonymous):

well i just substuited the x for c, lol i called my teacher and she told me what to do

OpenStudy (anonymous):

and then what did you do to get the answer?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

i believe you still cant get that answer

OpenStudy (anonymous):

There must be some other change in that question to get that answer... :)